### v1

# Did I Faithfully Say What I Thought? Bridging the Gap Between Neural Activity and Self-Explanations in Large Language Models

Abstract
Large Language Models (LLM) have demonstrated the capability of generating free text self Natural Language Explanation (self-NLE) to justify their answers. Despite their logical appearance, self-NLE do not necessarily reflect the LLM actual decision-making process, making such explanations unfaithful. While existing methods for measuring self-NLE faithfulness mostly rely on behavioral tests or computational block identification, none of them examines the neural activity underlying the model’s reasoning. This work introduces a novel flexible framework for quantitatively measuring the faithfulness of LLM-generated self-NLE by directly comparing the latter with interpretations of the model’s internal hidden states. The proposed framework is versatile and provides deep insights into self-NLE faithfulness by establishing a direct connection between self-NLE and model reasoning. This approach advances the understanding of self-NLE faithfulness and provides building blocks for generating more faithful self-NLE.
1 Introduction
Autoregressive Transformer-based Vaswani et al. (2017) Large Language Models (LLM) have demonstrated striking performance in a wide range of Natural Language Processing (NLP) tasks Chang et al. (2024). This achievement was largely attributable to the reasoning LLM abilities, enabling them to break down complex multi-step problems with Chain-of-Thought (CoT) Wei et al. (2024) and generate self free text Natural Language Explanation (self-NLE) to support their answers Wiegreffe and Marasovic ; Huang et al. (2023). Generating CoT consists in prompting the LLM to generate an explanation in an explain-then-predict fashion, intending to improve the model’s performance. Conversely, generating self-NLE follows a predict-then-explain setting, where explanation only has an explanatory purpose. Despite their logical and coherent appearance favoring trust in the model Han et al. (2023), LLM generated CoT and self-NLE turn out to not systematically reflect the actual underlying decision-making process of the model Agarwal et al. (2024), raising the need to accurately measure their faithfulness Jacovi and Goldberg (2020). This paper focuses on the case of measuring self-NLE faithfulness of LLM (i.e. in a predict-then-explain setting).
Numerous methods have been proposed to address this issue (see e.g. Lyu et al. (2024)). These approaches suffer from two main limitations: (1) they mostly perform behavioral tests and do not examine LLM internal reasoning processes Atanasova et al. (2023); Siegel et al. (2024, 2025); Matton et al. (2025), and (2) they identify the computational blocks that contribute to prediction and self-NLE without analyzing the actual neural activity underlying the prediction Wiegreffe et al. (2021); Parcalabescu and Frank (2024); Yeo et al. (2024).
To address these limitations, we propose NeuroFaith, a novel flexible framework for measuring the faithfulness of LLM self-NLE. The main contributions of this paper are as follows:
  1. 1.
    NeuroFaith enables to measure alignment between the LLM internal reasoning process and self-NLE by comparing the latter to a set of local neural interpretations decoding internal LLM hidden states at specific token indices and layers.
  2. 2.
    NeuroFaith is versatile and based on several components (location, circuit, interpreter), making it theoretically compatible with various linguistic tasks.
  3. 3.
    We instantiate NeuroFaith in 2-hop reasoning, and propose a thorough theoretical characterization of LLM self-NLE, according to their faithfulness and their correctness.
This paper is organized as follows: Section 2 recalls why measuring explanation faithfulness is important and presents how existing approaches measure the faithfulness of LLM self-NLE. Section 3 describes NeuroFaith in details. In Section 4 we instantiate NeuroFaith in 2-hop reasoning, and Section 5 discusses the conducted experiments.
2 Motivation & Related Work
2.1 Why Measure Explanation Faithfulness?
In their foundational paper Jacovi and Goldberg (2020) define model explanation faithfulness as "how accurately the explanation reflects the true reasoning process of the model", which is the definition generally adopted in the literature Lyu et al. (2024) and which we likewise follow throughout this work. Explanation faithfulness is commonly conflated with other explanation qualities such as readability and plausibility. This conflation can have serious consequences in critical domains, where explanations that appear plausible but lack faithfulness might lead end-users to over-rely on model predictions and make unfair Luo et al. (2022) or harmful Kayser et al. (2024) decisions.
This tension between plausibility and faithfulness has become more pronounced with recent advances in NLP Agarwal et al. (2024). Unlike their non-generative predecessors, modern LLM are trained to generate both answers and free text self-NLE that appear credible despite potentially containing persuasive hallucinations Rawte et al. (2023). The unstructured nature of these self-NLE pose greater challenges for faithfulness assessment compared to more constrained explainability techniques like attribution Wiegreffe et al. (2021) or counterfactual methods Madsen et al. (2024). As the use of LLM is expanding across diverse fields, the combination of their widespread adoption and the simplicity of generating self-NLE through prompting underpins the urgent to evaluate self-NLE faithfulness.
2.2 How to Measure Self-NLE Faithfulness of LLM?
Despite the critical importance of self-NLE faithfulness, there is no consensus on how to quantitatively measure it Lyu et al. (2024). Numerous approaches have been proposed that we group in 3 categories.
Counterfactual Interventions.
NLE faithfulness can be assessed through behavioral tests that measure how perturbations in the input text affect both predictions and self-NLE Atanasova et al. (2023). Counterfactual Intervention (CI) methods employ auxiliary models to generate counterfactual texts designed to change the LLM outcome. The LLM is then prompted to produce a self-NLE to justify its new prediction. The self-NLE is deemed faithful if it aligns with the specific CI that caused the prediction change.
These CI-based approaches mostly differ in two ways: how they measure consistency between the intervention and the resulting self-NLE Atanasova et al. (2023); Siegel et al. (2024, 2025), and the granularity of the CI methods, whether at token or concept level Matton et al. (2025). However, these approaches face several limitations: (1) the CI may not be solely responsible for the change in prediction, as the model might base its new prediction based on another part of the input text after intervention, (2) CI methods overlook the model’s underlying reasoning process, which departs from the commonly adopted definition of explanation faithfulness.
Attribution Agreement.
Another way to quantitatively measure self-NLE faithfulness is to compute post-hoc Attribution Agreement (AA) between the prediction and the NLE Parcalabescu and Frank (2024); Wiegreffe et al. (2021); Yeo et al. (2024). AA methods first compute attribution scores to interpret the model predictions and then measure how these scores correlate with other attribution scores computed to interpret related self-NLE. Higher correlation values indicate greater faithfulness in the model’s self-NLE.
AA approaches vary in their interpretation granularity (i.e., token or embedding level) and the post-hoc attribution method employed (gradient-based Sundararajan et al. (2017), SHAP Lundberg and Lee (2017) or activation patching Meng et al. (2022)). While AA methods assess whether the same LLM computational blocks were used when generating both the prediction and the self-NLE, they overlook the neural activity within these blocks, leaving the model’s inner reasoning process only partially treated.
Parametric Faithfulness.
The recent study by Tutek et al. (2025) emphasizes that CI methods actually measure contextual faithfulness by only perturbing the input text while keeping model parameters fixed. In contrast, it proposes to measure parametric faithfulness, which evaluates the faithfulness of self-NLE with respect to the model’s parameters. Their approach measures parametric faithfulness by assessing the effect of unlearning Liu et al. (2025) the parametric knowledge encoded in the reasoning steps of the self-NLE. The higher the change in prediction between the original model and the model having unlearned the reasoning steps, the more faithful the self-NLE. This approach inherently depends on the accuracy of the unlearning method employed, as either parameter localization or their erasure may be imprecise or inefficient for some target reasoning steps.
Most of the methods described above have been characterized as measuring self-consistency rather than true faithfulness Parcalabescu and Frank (2024), as they do not establish a direct link between the self-NLE and the model’s actual internal reasoning process. In the following, we propose a framework that improves upon these methods by explicitly examining the correspondence between self-NLE and interpretations of the model’s internal hidden states. However, we acknowledge our approach cannot perfectly capture the model’s complete inner reasoning process. Our approach therefore represents an advancement within the self-consistency paradigm by measuring consistency between neural internal activity and self-NLE. For convenience, in the following we keep on referring to faithfulness measurements.
3 NeuroFaith: A Framework for Measuring the Faithfulness of LLM self-NLE
This section introduces NeuroFaith, a flexible framework we propose to measure the faithfulness of LLM self-NLE by quantifying the alignment between the model’s internal reasoning processes and its self-NLE. This way, we categorize NeuroFaith as measuring neural activation faithfulness.
3.1 Overview
We consider a $L$-layer auto-regressive Transformer-based LLM $f$, a vocabulary space $\mathcal{V}$ and a text sequence of length $n$, $x=[t_{1},...,t_{n}]\in\mathcal{V}^{n}$. We denote $f(x)$ the answer and $e(x)$ the self-NLE related to $x$ produced by $f$. As pinpointed above, a faithful explanation should accurately reflect the model’s inner reasoning process. NeuroFaith aims to measure self-NLE faithfulness by comparing $e(x)$ to a set of local neural interpretations $i(x)=\left\{{i^{\ell}_{k}(x)}\right\}_{(k,\ell)\in\Omega}$, with $i^{\ell}_{k}(x)$ a decoded internal LLM hidden state at a specific token index $k$ and layer $\ell$ and $\Omega$ the set of couples (index/layer) whose hidden states are decoded. We define a set of local neural interpretations based on three components, detailed in turn in the following subsections:
  1. 1.
    A location $\mathcal{L}$, representing the granularity of the $f$ hidden states to be interpreted,
  2. 2.
    A circuit $\mathcal{C}$, defined as a set of indices and layers containing the hidden states of interest,
  3. 3.
    An interpreter $\mathcal{I}$, defined as an interpretability method designed to decode $f$ neural hidden states.
Each local neural interpretation $i^{\ell}_{k}(x)$ is compared to the self-NLE $e(x)$ to quantify its corresponding local faithfulness. The self-NLE is considered locally faithful if corroborated by the local neural interpretations. Finally, the global faithfulness is computed by aggregating the local faithfulness.
Figure 1: NeuroFaith overview depicted with location $\mathcal{L}=RS$, circuit $\mathcal{C}$ whose nodes are set as $\left\{{(12,1),(12,2)}\right\}$ and the interpretation method $\mathcal{I}$ set to either SAE or Selfie. The self-NLE justifies the model prediction ("Sweden") by stating that Ingmar Bergman directed the film Persona. NeuroFaith measures faithfulness by comparing this self-NLE with a set of interpretations that decode the model’s internal neural activity, represented as either concept-based interpretations or free text interpretations.
3.2 Location
To generate a set of local neural interpretations to be compared to the self-NLE, the first step is to define the specific neural architectural component, or location, $\mathcal{L}$ containing the hidden state that will be interpreted. Given an index $k$ and a layer $\ell$, a location $\mathcal{L}$ refers to a specific point within the model’s computational subgraph where neural activations can be analyzed and from which a local neural interpretation $i^{\ell}_{k}(x)$ is going to be generated.
Transformer-based generative LLM can be viewed as a stack of decoder computational blocks Radford et al. (2018). The information flow through a single layer and from a layer to another can be described with the following equations:
$h_{k}^{\ell}$ $=h_{k}^{\ell-1}+a_{k}^{\ell}+m_{k}^{\ell}$
$a_{k}^{\ell}$ $=\text{MHA}^{\ell}\left(h_{1}^{\ell-1},h_{2}^{\ell-1},\ldots,h_{k}^{\ell-1}\right)$
$m_{k}^{\ell}$ $=\text{MLP}^{\ell}(a_{k}^{\ell}+h_{k}^{\ell-1})$ (1)
where $h_{k}^{\ell}$ denotes the residual stream at index $k$ and layer $\ell$, $a_{k}^{\ell}$ the attention output of the multi-head attention operation $\text{MHA}^{\ell}$ and $m_{k}^{\ell}$ the output from the multi-layer perceptron $\text{MLP}^{\ell}$.
This characterization of the information flow through a Transformer block naturally highlights three possible locations, each offering different perspectives on the model’s information processing: the residual stream (RS) focusing on $h_{k}^{\ell}$, which serves as the main information pathway through the network; the multi-head attention (MHA) focusing on $a_{k}^{\ell}$, which captures token-to-token interactions; and the multi-layer perceptron (MLP) focusing on $m_{k}^{\ell}$ and performing non-linear transformations of the information Elhage et al. (2021); Rai et al. (2024). These locations offer a relevant granularity to decode the internal activity of an LLM, since focusing on neurons alone can rarely be done due to their polysemanticity Elhage et al. (2022). Thus, given a specific index/layer pair $(k,\ell)$ and following Equation 1, if $\mathcal{L}=\texttt{RS}$, $i^{\ell}_{k}(x)$ can be defined as the local neural interpretation of the hidden state $h^{\ell}_{k}$ on the residual stream.
[…]
3.3 Circuit
We restrict the interpretation of the hidden states to computational subgraphs that most significantly influences $f$ prediction by focusing on circuits Elhage et al. (2021). A circuit represents a sparse oriented subgraph of a neural network having an interpretable functional role, where nodes correspond to computational units (e.g. neuron, attention head, RS hidden state) and edges represent computation paths Wang et al. (2023); Räuker et al. (2023).
Formally, a circuit can be defined as $\mathcal{C}=(\mathcal{C}_{V},\mathcal{C}_{E})$ where $\mathcal{C}_{V}$ represents the set of computational nodes and $\mathcal{C}_{E}$ represents the set of directed edges between nodes where the information flows. We assume in the following that each computational node in circuit $\mathcal{C}$ is defined by a specific pair of coordinates: an index and a layer. The computational units at these nodes are specified by location $\mathcal{L}$. Thus, $\mathcal{C}_{V}$ consists of all unique index-layer pairs relevant to the reasoning process. In this way, $\mathcal{C}$ is restricted to have a particular form, being defined by its coordinates in terms of token index and model’s layer, and being unable to have computational units finer than the granularity imposed by the choice of location $\mathcal{L}$. Generating a set of local neural interpretations thus requires defining a circuit $\mathcal{C}$ based on computational units dependent on a location $\mathcal{L}$. Circuits can be obtained in different ways.
Task Specific Recurrent Circuits.
A vast body of research has manually investigated task-specific circuits in Transformer-based LLM Bereska and Gavves (2024), such as entity resolution Wang et al. (2023), induction heads Olsson et al. (2022) or multi-hop reasoning Biran et al. (2024). Analyzing the faithfulness of an NLE for such tasks would consists in directly drawing inspiration from these established circuits to define $\mathcal{C}$.
Automated Circuit Discovery.
Alternatively, circuits can be obtained through automated methods, including techniques based on activation patching Meng et al. (2022); Syed et al. (2024); Conmy et al. (2023), backward attribution Ferrando et al. (2022); Ferrando and Voita (2024) or Transcoders Dunefsky et al. (2024); Ameisen et al. (2025). These automated approaches help address scenarios where circuits may not be well-documented in the literature.
A set of local neural interpretations is thus defined by a location $\mathcal{L}$ and based on a circuit $\mathcal{C}$, where interpretations are generated for each node (index/layer pairs) in the circuit, with $\Omega=\mathcal{C}_{V}$.
3.4 Post-hoc Hidden State Interpreter
To generate a set of interpretations, the final step is to select an interpretability method $\mathcal{I}$ to decode $f$ hidden states on location $\mathcal{L}$ across circuit $\mathcal{C}$. We focus on interpretations expressed either as concepts or in the form of free text. These approaches provide insights into the internal neural activity of a language model, offering a valuable approximation of $f$ reasoning process for measuring the faithfulness of its self-NLE.
Concept-based Interpretation.
Recent research has shown that interpretable high-level features, referred to as concepts, appear to be linearly encoded within LLM latent space Elhage et al. (2022); Park et al. (2024). This computational understanding of concepts aligns with definitions from cognitive science Ruiz Luyten and van der Schaar (2024), where concepts are considered as mental entities essential to thought, enabling information integration and categorization Goguen (2005). This alignment makes concepts an ideal granularity level for model interpretation. Concept-based interpretations can be either obtained in a supervised or unsupervised manner, depending on the availability of texts with labeled concepts.
We make the general assumption that no such set of labeled concepts is available. Generating concept-based interpretations can be done by applying dictionary learning techniques Tošić and Frossard (2011) to project hidden states into a concept space. The semantics of the concepts are then determined once the concept dictionary has been learned. A popular approach for learning such dictionaries is to train sparse autoencoders (SAEHuben et al. (2023); Bricken et al. (2023) on the hidden states of $f$. Sparse concept activations are retrieved in the SAE hidden space to construct the concept-based activations as an interpretation.
This way, if $\mathcal{I}=\texttt{SAE}$, the interpretation at index $k$ and layer $\ell$ is defined as $i^{\ell}_{k}(x)\in\mathbb{R}^{d}$, where $d$ is the dimension of the concept dictionary and each element represents the activation of a specific concept.
Free Text Interpretation.
Another way of interpreting $f$ hidden states is to generate free text interpretations. Two recent and similar approaches, namely Selfie Chen et al. (2024) and Patchscopes Ghandeharioun et al. (2024), perform such interpretation generation by leveraging LLM decoding capabilities to produce natural language descriptions of hidden states. The interpretation is done by replacing the embedding of a placeholder token during a forward pass with the hidden state to be interpreted. The prompt of the interpretation forward pass is constructed to enhance free text interpretation expressivity.
This way, if $\mathcal{I}=\texttt{Selfie}$, the interpretation at index $k$ and layer $\ell$ is projected into the vocabulary space and defined as a text sequence $i^{\ell}_{k}(x)\in\mathcal{V}^{D}$, with $D$ the maximum length of the generated interpretation.
Finally, given an input text $x$, a set of local neural interpretations $i(x)$ is derived by applying the chosen interpreter $\mathcal{I}$ to a set of hidden states defined by their location $\mathcal{L}$ across a circuit $\mathcal{C}$, where each interpretation is generated with respect to an index/layer pair. The set of interpretations can then be formally defined as $i(x)=\bigcup\limits_{k,\ell\in\mathcal{C}_{V}}i_{k}^{\ell}(x)$.
3.5 Measuring Faithfulness
We assess the faithfulness of the self-NLE $e(x)$ by comparing it with the set of interpretations $i(x)$. Our aim is to measure the extent to which $e(x)$ can be corroborated or contradicted by $i(x)$, depending on the format (concept-based or free-text) of the latter. In the following, for simplicity of notation, we denote $F(x,e,i)$ the global faithfulness measure of $e(x)$ with respect to $i(x)$ and $F(x,e,i^{\ell}_{k})$ the local faithfulness measure of $e(x)$ with respect to a specific interpretation $i_{k}^{\ell}(x)$ on an index/layer pair $(k,\ell)$.
Local Faithfulness.
For $\mathcal{I}=\texttt{SAE}$, each interpretation in $i(x)$ is a set of vectors of concept activations with dimensions potentially exceeding a million Templeton et al. (2024), making direct comparison challenging. For a given interpretation $i^{\ell}_{k}(x)$, one way to simplify this comparison is to retrieve the concept labels with the highest activations. We formally define this simplification as applying the transformation $g(i^{\ell}_{k}(x))$ with $g:\mathbb{R}^{d}\rightarrow\mathcal{V}^{d^{\prime}}$ and $d^{\prime}\ll d$. The local faithfulness measure then evaluates whether the $d^{\prime}$ most salient concepts from $g(i^{\ell}_{k}(x))$ are present in $e(x)$.
For $\mathcal{I}=\texttt{Selfie}$, each interpretation in $i(x)$ consists of free-text interpretations. In this case, an interpretation $i^{\ell}_{k}(x)$ can be compared to $e(x)$ by assessing whether one logically corroborates the other. This comparison can be done either under an LLM-as-a-Judge settings Gu et al. (2024) or through human evaluation. We give a detailed example of how this comparison can be carried out in the case of 2-hop reasoning in Section 4.
Global Faithfulness.
We define the global faithfulness measure of $e(x)$ with respect to $i(x)$ as an aggregation of all local faithfulness measures of $e(x)$ across the nodes $\mathcal{C}_{V}$ of circuit $\mathcal{C}$:
$F(x,e,i)=\underset{(k,\ell)\in\mathcal{C}_{V}}{\bigodot}\left\{F(x,e,i^{\ell}_{k})\right\}$ (2)
[…]
4 The Case of 2-Hop Reasoning
In this section, we demonstrate why 2-hop reasoning is convenient for measuring self-NLE faithfulness and illustrate how self-NLE faithfulness can be assessed in this context by comparison with interpretations at the bridge object level. We instantiate NeuroFaith in the case of 2-hop reasoning, and precisely characterize self-NLE with respect to both its faithfulness and its correctness.
4.1 Task Description
Multi-hop reasoning is a complex cognitive task in NLP that requires connecting a sequence of information to reach a conclusion Mavi et al. (2024). It consists in several single-hop operations Trivedi et al. (2022), which can individually be defined as triplets $(o_{1},r,o_{2})$ where $o_{1}$ is a source object, $r$ is a relation and $o_{2}$ is a target object. For example, the 2-hop reasoning statement "The country of origin of the movie maker that directed the movie Persona is Sweden" requires sequentially solving the two single-hop operations: $(o_{1},r_{1},o_{2})$ and $(o_{2},r_{2},o_{3})$ where $o_{1}$ is "Persona" (the movie), $r_{1}$ is "movie direction", $o_{2}$ is "Ingmar Bergman", $r_{2}$ is "country of origin" and the final answer $o_{3}$ is "Sweden". In this example, $o_{2}$ ("Ingmar Bergman") serves as the bridge object, connecting two reasoning operations and representing the intermediate step necessary to generate the correct final answer.
[…]
Definition 1: Answer Correctness If $\bm{\widehat{o_{3}}=o_{3}}$, the final answer is correct.
In the following we focus on the case of 2-hop reasoning and discuss why it is particularly well-suited for faithfulness measurement.
4.2 Measuring the Faithfulness of 2-Hop Self-NLE
The 2-hop reasoning task provides an ideal framework for measuring self-NLE faithfulness due to its single, explicit intermediate step. Unlike more complex reasoning tasks, 2-hop reasoning exposes a single, clear bridge object ($o_{2}$) that connects two reasoning operations, making it easier to verify if a $f$ self-NLE aligns with its internal reasoning process. Beyond 2-hop reasoning (with 3+ hops), measuring faithfulness becomes significantly more challenging as the number of bridge objects increases. Multiple bridge objects are likely to introduce cascading effects where errors at early hops influence subsequent reasoning steps, making it difficult to determine whether later steps are faithful to $f$ internal representations.
We consider a self-NLE $e(x)$ generated in a predict-then-explain fashion for input $x$ and its corresponding $f$ answer $\widehat{o_{3}}$. Two reasoning chains are derived from $e(x)$: $(o_{1},r_{1},\widehat{o_{2}})$ and $(\widehat{o_{2}},r_{2},\widehat{o_{3}})$, where $\widehat{o_{2}}$ represents the bridge object mentioned in the self-NLE $e(x)$. Besides, we assume we have a set of local neural interpretations $i(x)=\left\{{i^{\ell}_{k}(x)}\right\}_{(k,\ell)}$ previously obtained by appling NeuroFaith. We argue that, given a token index $k$ and a layer $\ell$, $e(x)$ is locally faithful if $\widehat{o_{2}}$ is contained in the $i^{\ell}_{k}(x)$ interpretation of the $f$ hidden state, as it then aligns with the model’s internal reasoning process.
Definition 2: Self-NLE Local Faithfulness If $\bm{\widehat{o_{2}}\in i^{\ell}_{k}(x)}$, $\bm{F(x,e,i^{\ell}_{k})=1}$ and the self-NLE is considered locally faithful with respect to $k$ and $\ell$.
Conversely, if $\widehat{o_{2}}\notin i^{\ell}_{k}(x)$, $F(x,e,i^{\ell}_{k})=0$ and we consider the self-NLE locally unfaithful with respect to $k$ and $\ell$, as it is misaligned with the $f$ internal reasoning process.
4.3 Measuring the Correctness of Self-NLE, Interpretation and Latent 2-hop Reasoning
In 2-hop reasoning, focusing on the bridge object enables to bring out an additional dimension of characterization of $e(x)$ beyond faithfulness, regarding the expected ground truth bridge object $o_{2}$. The self-NLE $e(x)$ is said to be correct if its corresponding bridge object is the same as the expected ground truth one.
Definition 3: Self-NLE Correctness If $\bm{\widehat{o_{2}}=o_{2}}$, the self-NLE is correct.
In the same way, we can define the correctness of the interpretation $i(x)$ through comparison with $o_{2}$.
Definition 4: Interpretation Correctness If $\exists(k,l)\in\mathcal{C}_{V}$ such that $o_{2}\in i^{\ell}_{k}(x)$, the interpretation $i(x)$ is correct.
We can then define the correctness of the 2-hop reasoning task, based on answer correctness and local neural interpretation correctness.
Definition 5: Latent 2-hop Correctness If $\bm{\widehat{o_{3}}=o_{3}}$ and if $\exists(k,l)\in\mathcal{C_{V}}$ such that $o_{2}\in i^{\ell}_{k}(x)$, $f$ correctly performs latent 2-hop reasoning.
This characterization of latent 2-hop reasoning is consistent with the one already mentioned in previous work Yang et al. (2024); Biran et al. (2024).
4.4 Highlighting the Tension Between Correctness and Faithfulness of 2-hop Self-NLE
The correctness of the prediction combined with both the faithfulness and the correctness of the self-NLE enables to precisely characterize $e(x)$. In this subsection, we focus on five disjoint cases of interest to characterize the behavior of $f$ with respect to $e(x)$ when the prediction is correct (i.e. $\widehat{o_{3}}=o_{3}$). They are illustrated with examples of cases above in Appendix A.1.
  • ($C_{1}$) $f$ as a reliable oracle. If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=1$ and $\widehat{o_{2}}=o_{2}$, the final prediction is correct, the self-NLE is correct, and faithful at the same time. This case is the most favorable, and sometimes enables knowledge transfer based on $f$ self-NLE.
  • ($C_{2}$) $f$ being biased. If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=1$ and $\widehat{o_{2}}\neq o_{2}$, the final prediction is correct, the self-NLE is faithful but incorrect, suggesting a potential bias or confounder (possibly contained in $i(x)$) making $f$ predict accurately but for the wrong reason.
  • ($C_{3}$) $f$ as an explainer parrot. If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$ and $\widehat{o_{2}}=o_{2}$, the prediction is correct and the self-NLE identifies the ground truth bridge object while being globally unfaithful to $f$ internal reasoning. This case suggests that $f$ does not perform latent 2-hop reasoning properly and merely replicates the expected explanation once the answer is known.
  • ($C_{4}$) $f$ being deceptive. If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$, $\widehat{o_{2}}\neq o_{2}$ and $o_{2}\in i^{\ell}_{k}(x)$, $f$ appears to correctly perform latent reasoning but does not faithfully express this in its self-NLE, this way doing deceptive reasoning Greenblatt et al. (2024). This case is expected to be rare, otherwise highlighting a case where $f$ is not honest in its self-NLE while knowing the ground truth bridge object, raising a problem in $f$ alignment.
  • ($C_{5}$) $f$ doing shortcut learning. If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\mathcal{C_{V}},o_{2}\notin i^{\ell}_{k}(x)$, the prediction is correct while $f$ seems to neither correctly perform latent reasoning nor generate a correct and faithful self-NLE, suggesting a potential case of shortcut learning Geirhos et al. (2020). This case is different from $C_{2}$ since shortcut learning implies correctly predicting $o_{3}$ directly based on $o_{1}$, whereas when $f$ is biased, the $o_{3}$ prediction is based on an incorrect bridge object $o_{2}$.
5 Experimental Settings
This section presents the experimental study applying NeuroFaith to the case of 2-hop reasoning with two decoder-only LLM of different sizes.
5.1 Datasets and Models
We apply NeuroFaith on the Wikidata-2-hop dataset Biran et al. (2024). Following the notations introduced in Section 4, each input text is formatted as $x=(o_{1},r_{1},{\blacktriangle{}},r_{2},\bullet)$. We evaluate Gemma-2-2B and Gemma-2-9B Riviere et al. (2024) in their instruction-tuned versions, using the following backbones from the transformers library Wolf et al. (2020): gemma-2-2B-it and gemma-2-9B-it.
5.2 Framework Instantiation
Our experimental process follows 5 steps:
  1. 1.
    Answer Generation. We first prompt $f$ to generate an answer $\widehat{o_{3}}$, given an input text $x=(o_{1},r_{1},{\blacktriangle{}},r_{2},\bullet)$ using instruction: "Complete the following text" followed by $x$. We then compute the final 2-hop reasoning correctness (1 if $\widehat{o_{3}}=o_{3}$, 0 otherwise).
  2. 2.
    Self-NLE Generation. Once the answer $\widehat{o_{3}}$ has been generated, we ask $f$ to generate a self-NLE $e(x)$ in a predict-then-explain fashion with the following prompt: "Complete the following text and give a short justification" followed by $(o_{1},r_{1},{\blacktriangle{}},r_{2},\widehat{o_{3}})$.
  3. 3.
    Bridge Object Extraction. We extract the bridge object $\widehat{o_{2}}$ from $e(x)$ using the model Qwen3-32B with the prompt: "What is the entity linking $o_{1}$ and $\widehat{o_{3}}$ in the following text?" followed by $e(x)$. We then compute the self-NLE correctness (1 if $\widehat{o_{2}}=o_{2}$, 0 otherwise).
  4. 4.
    Hidden States Interpretation. We use Selfie Chen et al. (2024) as interpreter $\mathcal{I}$ to decode the $f$ hidden states of interest and generate a set of local neural interpretations $i(x)$. We focus on the residual stream ($\mathcal{L}=\texttt{RS}$) and define the circuit $\mathcal{C}$ based on recent work Biran et al. (2024); Yang et al. (2024), which identifies the latent first hop operation in the early layers (when processing the last source token corresponding $o_{1}$). We therefore set the index token $k$ as the last one related to $o_{1}$ and $\ell\in\left\{{5,6,7,8,9,10,11}\right\}$ for gemma-2-2B-it and $\ell\in\left\{{8,9,10,11,12,13,14}\right\}$ for gemma-2-9B-it.
  5. 5.
    Self-NLE Faithfulness. We finally compare the set of interpretations $i(x)$ to $\widehat{o_{3}}$ to assess the faithfulness of $e(x)$. Given an index $k$ and a layer $\ell$, the local faithfulness $F(x,e,i_{k}^{\ell})$ is measured following the definition introduced in Section 4.2. The global faithfulness $F(x,e,i)$ is computed in the soft way with the $\max$ aggregation function.
We finally evaluate the 2-hop reasoning accuracy (ACC), the self-NLE correctness (NLE-c), the interpretation correctness (I-c) and the self-NLE faithfulness (NLE-f) on the entire dataset. We also consider the five cases introduced in Section 4.4, referred to as $C_{p}$ for a specific case $p$.
Metric Gemma-2-2B Gemma-2-9B
ACC (%) 24.7 28.9*
NLE-f (%) 22.8 23.6*
$\widehat{o_{3}}=o_{3}$ 25.2 29.2
$\widehat{o_{3}}\neq o_{3}$ 22.1 15.7
NLE-c (%) 62.3*** 52.9
$\widehat{o_{3}}=o_{3}$ 91.3 57.3
$\widehat{o_{3}}\neq o_{3}$ 36.2 33.7
I-c (%) 17.8 21.6**
$\widehat{o_{3}}=o_{3}$ 23.3 28.6
$\widehat{o_{3}}\neq o_{3}$ 15.5 11.8
Table 1: NeuroFaith evaluation per model size on the Wikidata-2-hop dataset. Results are presented in percentages. The best results are highlighted in bold. With $p$ as the $p$-value of the $t$-test comparing Gemma-2-2B and Gemma-2-9B all prediction types combined, *$p<10$%, **$p<5$%, ***$p<1$%.
Metric (%) Gemma-2-2B Gemma-2-9B
$C_{1}$
$f$ as a reliable oracle
18.7 28.5***
$C_{2}$
$f$ being biased
6.5*** 0.7
$C_{3}$
$f$ as an explainer parrot
37.9 40.4*
$C_{4}$
$f$ being deceptive
0 0
$C_{5}$
$f$ doing shortcut learning
8.7 34.4***
Table 2: Representativeness of cases of interest obtained from NeuroFaith by model size on the Wikidata-2-hop dataset. Results are presented in percentages. With $p$ as the $p$-value of the $t$-test comparing Gemma-2-2B and Gemma-2-9B, *$p<10$%, **$p<5$%, ***$p<1$%.
5.3 Results
Global Results.
Table 1 shows that overall, both models are generally not very faithful, although in most cases they mention the bridge object in their self-NLE. Gemma-2-9B achieves the best results in terms of final task accuracy (28.9%) as compared to Gemma-2-2B (24.7%). It generates slightly more faithful self-NLE but significantly less correct self-NLE. It tends to correctly achieve the first-hop in the latent space more often. The status of the final prediction (accurate or not) has a significant impact on the faithfulness and correctness of the self-NLE and the correctness of the interpretation, as shown by the rows detailing the values depending on whether $\widehat{o_{3}}=o_{3}$ or not . Overall, self-NLE are more faithful and correct, and interpretations are more correct when the final prediction is correct.
Self-NLE Characterization Analysis.
Table 2 presents the distribution of the case types defined in Section 4.4 among correctly predicted texts (i.e. $\widehat{o_{3}}=o_{3}$). Gemma-2-9B proves to be a more reliable oracle ($C_{1}$) than Gemma-2-2B, while also being less biased ($C_{2}$), producing more self-NLE that are faithful but incorrect. Both models function as explainer parrots at similar rates (approximately 40%), thus often generating self-NLE that are correct but unfaithful. This high value seems to highlight a flagrant case of misalignment. Finally, Gemma-2-9B engages in shortcut learning significantly more often than Gemma-2-2B, correctly producing the final answer without evidence of the expected bridge object in its latent space or self-NLE. This shortcut learning behavior represents the second most common scenario when Gemma-2-9B correctly determines the expected answer. These values must be taken with caution, as they may be due to the choice of a circuit too small to decode the model’s internal states.
6 Discussion
In this paper, we focus on generating free text interpretations for faithfulness assessment. Future work could explore additional NeuroFaith implementations, such as using Sparse Autoencoders (SAE) to generate concept-based interpretations. Additionally, our detailed characterization focused exclusively on cases with correct final predictions ($\widehat{o_{3}}=o_{3}$). Developing similar characterizations when models make incorrect predictions would provide valuable insights. Our current study calculated faithfulness of free text self-NLE only in predict-then-explain contexts. A promising extension would be to compare the faithfulness of self-NLE with the performance gains they can induce Bhan et al. (2024). We believe that NeuroFaith could also lead to valuable results by measuring CoT faithfulness and being compared to other studies in this direction Lanham et al. (2023); Turpin et al. (2023). Examining the faithfulness of alternative explanation methods such as counterfactuals Madsen et al. (2024) or concept-based approaches Bhan et al. (2025) through comparison with the model’s internal neural activity is another promising research direction.
7 Conclusion
In this paper we introduced NeuroFaith, a new flexible framework that integrates with diverse interpretability approaches for decoding LLM internal neural processes and measuring self-NLE faithfulness that complements existing literature. NeuroFaith establishes a stronger connection between self-NLE and the model’s internal reasoning processes, aligning more closely with the fundamental definition of faithfulness. This work enhances our understanding of self-NLE faithfulness and creates a foundation for developing more faithful self-explanations in LLM.
Appendix A Appendix
A.1 Examples of Characterizations of Cases of Interest
Figure 2: Example from the Wikidata-2-hop dataset where model $f$ acts as a reliable oracle. The model correctly answers "The United States" and provides a self-NLE referencing the expected bridge object "Pierre Omidyar". This same bridge object appears in a free text interpretation at token position 9 and layer 8, demonstrating that the answer is correct while the self-NLE is simultaneously correct and faithful.
Figure 3: Example from the Wikidata-2-hop dataset where model $f$ acts as being biased. The model correctly answers "Japan" and provides a self-NLE referencing the bridge object "Square Enix" whereas the expected bridge object is "Nintendo". This same bridge object appears in a free text interpretation at token position 10 and layer 9, demonstrating that the answer is correct while the self-NLE is simultaneously wrong and faithful.
Figure 4: Example from the Wikidata-2-hop dataset where model $f$ seems to act as an explainer parrot. The model correctly answers "Germany" and provides a self-NLE referencing the correct bridge object "Wim Wenders". This expected bridge object does not appear in any free text interpretation in the circuit, demonstrating that the answer is correct while the self-NLE is simultaneously correct and unfaithful.
Figure 5: Example from the Wikidata-2-hop dataset where model $f$ seems to perform shortcut learning. The model correctly answers "France" and provides a self-NLE referencing the wrong bridge object "Alain Resnais". Neither this wrong bridge object nor the expected one appear in the set of free text interpretations across the circuit, demonstrating that the answer is correct while the self-NLE is simultaneously incorrect and unfaithful.

### v2

@@
 1 Introduction
-Autoregressive Transformer-based Vaswani et al.
-(2017) Large Language Models (LLM) have demonstrated striking performance in a wide range of Natural Language Processing (NLP) tasks Chang et al.
-(2024).
-This achievement was largely attributable to the reasoning LLM abilities, enabling them to break down complex multi-step problems with Chain-of-Thought (CoT) Wei et al.
-(2024) and generate self free text Natural Language Explanation (self-NLE) to support their answers Wiegreffe and Marasovic ; Huang et al.
-(2023).
+Autoregressive Transformer-based (Vaswani et al., 2017) Large Language Models (LLM) have demonstrated striking performance in a wide range of Natural Language Processing (NLP) tasks (Chang et al., 2024).
+This achievement was largely attributable to the reasoning LLM abilities, enabling them to break down complex multi-step problems with Chain-of-Thought (CoT) (Wei et al., 2024) and generate self free text Natural Language Explanation (self-NLE) to support their answers (Wiegreffe and Marasovic, ; Huang et al., 2023).
 Generating CoT consists in prompting the LLM to generate an explanation in an explain-then-predict fashion, intending to improve the model’s performance.
 Conversely, generating self-NLE follows a predict-then-explain setting, where explanation only has an explanatory purpose.
-Despite their logical and coherent appearance favoring trust in the model Han et al.
-(2023), LLM generated CoT and self-NLE turn out to not systematically reflect the actual underlying decision-making process of the model Agarwal et al.
-(2024), raising the need to accurately measure their faithfulness Jacovi and Goldberg (2020).
+Despite their logical and coherent appearance favoring trust in the model (Han et al., 2023), LLM generated CoT and self-NLE turn out to not systematically reflect the actual underlying decision-making process of the model (Agarwal et al., 2024), raising the need to accurately measure their faithfulness (Jacovi and Goldberg, 2020).
 This paper focuses on the case of measuring self-NLE faithfulness of LLM (i.e.
@@
 (2024)).
-These approaches suffer from two main limitations: (1) they mostly perform behavioral tests and do not examine LLM internal reasoning processes Atanasova et al.
-(2023); Siegel et al.
-(2024, 2025); Matton et al.
-(2025), and (2) they identify the computational blocks that contribute to prediction and self-NLE without analyzing the actual neural activity underlying the prediction Wiegreffe et al.
-(2021); Parcalabescu and Frank (2024); Yeo et al.
-(2024).
+These approaches suffer from two main limitations: (1) they mostly perform behavioral tests and do not examine LLM internal reasoning processes (Atanasova et al., 2023; Siegel et al., 2024, 2025; Matton et al., 2025), and (2) they identify the computational blocks that contribute to prediction and self-NLE without analyzing the actual neural activity underlying the prediction (Wiegreffe et al., 2021; Parcalabescu and Frank, 2024; Yeo et al., 2024).
 To address these limitations, we propose NeuroFaith, a novel flexible framework for measuring the faithfulness of LLM self-NLE. The main contributions of this paper are as follows:
@@
 2.1 Why Measure Explanation Faithfulness?
-In their foundational paper Jacovi and Goldberg (2020) define model explanation faithfulness as "how accurately the explanation reflects the true reasoning process of the model", which is the definition generally adopted in the literature Lyu et al.
-(2024) and which we likewise follow throughout this work.
+In their foundational paper Jacovi and Goldberg (2020) define model explanation faithfulness as "how accurately the explanation reflects the true reasoning process of the model", which is the definition generally adopted in the literature (Lyu et al., 2024) and which we likewise follow throughout this work.
 Explanation faithfulness is commonly conflated with other explanation qualities such as readability and plausibility.
-This conflation can have serious consequences in critical domains, where explanations that appear plausible but lack faithfulness might lead end-users to over-rely on model predictions and make unfair Luo et al.
-(2022) or harmful Kayser et al.
-(2024) decisions.
-This tension between plausibility and faithfulness has become more pronounced with recent advances in NLP Agarwal et al.
-(2024).
-Unlike their non-generative predecessors, modern LLM are trained to generate both answers and free text self-NLE that appear credible despite potentially containing persuasive hallucinations Rawte et al.
-(2023).
-The unstructured nature of these self-NLE pose greater challenges for faithfulness assessment compared to more constrained explainability techniques like attribution Wiegreffe et al.
-(2021) or counterfactual methods Madsen et al.
-(2024).
+This conflation can have serious consequences in critical domains, where explanations that appear plausible but lack faithfulness might lead end-users to over-rely on model predictions and make unfair (Luo et al., 2022) or harmful (Kayser et al., 2024) decisions.
+This tension between plausibility and faithfulness has become more pronounced with recent advances in NLP (Agarwal et al., 2024).
+Unlike their non-generative predecessors, modern LLM are trained to generate both answers and free text self-NLE that appear credible despite potentially containing persuasive hallucinations (Rawte et al., 2023).
+The unstructured nature of these self-NLE pose greater challenges for faithfulness assessment compared to more constrained explainability techniques like attribution (Wiegreffe et al., 2021) or counterfactual methods (Madsen et al., 2024).
 As the use of LLM is expanding across diverse fields, the combination of their widespread adoption and the simplicity of generating self-NLE through prompting underpins the urgent to evaluate self-NLE faithfulness.
 2.2 How to Measure Self-NLE Faithfulness of LLM?
-Despite the critical importance of self-NLE faithfulness, there is no consensus on how to quantitatively measure it Lyu et al.
-(2024).
+Despite the critical importance of self-NLE faithfulness, there is no consensus on how to quantitatively measure it (Lyu et al., 2024).
 Numerous approaches have been proposed that we group in 3 categories.
 Counterfactual Interventions.
-NLE faithfulness can be assessed through behavioral tests that measure how perturbations in the input text affect both predictions and self-NLE Atanasova et al.
-(2023).
+NLE faithfulness can be assessed through behavioral tests that measure how perturbations in the input text affect both predictions and self-NLE (Atanasova et al., 2023).
 Counterfactual Intervention (CI) methods employ auxiliary models to generate counterfactual texts designed to change the LLM outcome.
@@
 The self-NLE is deemed faithful if it aligns with the specific CI that caused the prediction change.
-These CI-based approaches mostly differ in two ways: how they measure consistency between the intervention and the resulting self-NLE Atanasova et al.
-(2023); Siegel et al.
-(2024, 2025), and the granularity of the CI methods, whether at token or concept level Matton et al.
-(2025).
+These CI-based approaches mostly differ in two ways: how they measure consistency between the intervention and the resulting self-NLE (Atanasova et al., 2023; Siegel et al., 2024, 2025), and the granularity of the CI methods, whether at token or concept level (Matton et al., 2025).
 However, these approaches face several limitations: (1) the CI may not be solely responsible for the change in prediction, as the model might base its new prediction based on another part of the input text after intervention, (2) CI methods overlook the model’s underlying reasoning process, which departs from the commonly adopted definition of explanation faithfulness.
 Attribution Agreement.
-Another way to quantitatively measure self-NLE faithfulness is to compute post-hoc Attribution Agreement (AA) between the prediction and the NLE Parcalabescu and Frank (2024); Wiegreffe et al.
-(2021); Yeo et al.
-(2024).
+Another way to quantitatively measure self-NLE faithfulness is to compute post-hoc Attribution Agreement (AA) between the prediction and the NLE (Parcalabescu and Frank, 2024; Wiegreffe et al., 2021; Yeo et al., 2024).
 AA methods first compute attribution scores to interpret the model predictions and then measure how these scores correlate with other attribution scores computed to interpret related self-NLE.
 Higher correlation values indicate greater faithfulness in the model’s self-NLE.
-AA approaches vary in their interpretation granularity (i.e., token or embedding level) and the post-hoc attribution method employed (gradient-based Sundararajan et al.
-(2017), SHAP Lundberg and Lee (2017) or activation patching Meng et al.
-(2022)).
+AA approaches vary in their interpretation granularity (i.e., token or embedding level) and the post-hoc attribution method employed (gradient-based (Sundararajan et al., 2017), SHAP (Lundberg and Lee, 2017) or activation patching (Meng et al., 2022)).
 While AA methods assess whether the same LLM computational blocks were used when generating both the prediction and the self-NLE, they overlook the neural activity within these blocks, leaving the model’s inner reasoning process only partially treated.
@@
 In contrast, it proposes to measure parametric faithfulness, which evaluates the faithfulness of self-NLE with respect to the model’s parameters.
-Their approach measures parametric faithfulness by assessing the effect of unlearning Liu et al.
-(2025) the parametric knowledge encoded in the reasoning steps of the self-NLE.
+Their approach measures parametric faithfulness by assessing the effect of unlearning (Liu et al., 2025) the parametric knowledge encoded in the reasoning steps of the self-NLE.
 The higher the change in prediction between the original model and the model having unlearned the reasoning steps, the more faithful the self-NLE.
 This approach inherently depends on the accuracy of the unlearning method employed, as either parameter localization or their erasure may be imprecise or inefficient for some target reasoning steps.
-Most of the methods described above have been characterized as measuring self-consistency rather than true faithfulness Parcalabescu and Frank (2024), as they do not establish a direct link between the self-NLE and the model’s actual internal reasoning process.
+Most of the methods described above have been characterized as measuring self-consistency rather than true faithfulness (Parcalabescu and Frank, 2024), as they do not establish a direct link between the self-NLE and the model’s actual internal reasoning process.
 In the following, we propose a framework that improves upon these methods by explicitly examining the correspondence between self-NLE and interpretations of the model’s internal hidden states.
@@
 Given an index $k$ and a layer $\ell$, a location $\mathcal{L}$ refers to a specific point within the model’s computational subgraph where neural activations can be analyzed and from which a local neural interpretation $i^{\ell}_{k}(x)$ is going to be generated.
-Transformer-based generative LLM can be viewed as a stack of decoder computational blocks Radford et al.
-(2018).
+Transformer-based generative LLM can be viewed as a stack of decoder computational blocks (Radford et al., 2018).
 The information flow through a single layer and from a layer to another can be described with the following equations:
@@
 where $h_{k}^{\ell}$ denotes the residual stream at index $k$ and layer $\ell$, $a_{k}^{\ell}$ the attention output of the multi-head attention operation $\text{MHA}^{\ell}$ and $m_{k}^{\ell}$ the output from the multi-layer perceptron $\text{MLP}^{\ell}$.
-This characterization of the information flow through a Transformer block naturally highlights three possible locations, each offering different perspectives on the model’s information processing: the residual stream (RS) focusing on $h_{k}^{\ell}$, which serves as the main information pathway through the network; the multi-head attention (MHA) focusing on $a_{k}^{\ell}$, which captures token-to-token interactions; and the multi-layer perceptron (MLP) focusing on $m_{k}^{\ell}$ and performing non-linear transformations of the information Elhage et al.
-(2021); Rai et al.
-(2024).
-These locations offer a relevant granularity to decode the internal activity of an LLM, since focusing on neurons alone can rarely be done due to their polysemanticity Elhage et al.
-(2022).
+This characterization of the information flow through a Transformer block naturally highlights three possible locations, each offering different perspectives on the model’s information processing: the residual stream (RS) focusing on $h_{k}^{\ell}$, which serves as the main information pathway through the network; the multi-head attention (MHA) focusing on $a_{k}^{\ell}$, which captures token-to-token interactions; and the multi-layer perceptron (MLP) focusing on $m_{k}^{\ell}$ and performing non-linear transformations of the information (Elhage et al., 2021; Rai et al., 2024).
+These locations offer a relevant granularity to decode the internal activity of an LLM, since focusing on neurons alone can rarely be done due to their polysemanticity (Elhage et al., 2022).
 Thus, given a specific index/layer pair $(k,\ell)$ and following Equation 1, if $\mathcal{L}=\texttt{RS}$, $i^{\ell}_{k}(x)$ can be defined as the local neural interpretation of the hidden state $h^{\ell}_{k}$ on the residual stream.
@@
 3.3 Circuit
-We restrict the interpretation of the hidden states to computational subgraphs that most significantly influences $f$ prediction by focusing on circuits Elhage et al.
-(2021).
+We restrict the interpretation of the hidden states to computational subgraphs that most significantly influences $f$ prediction by focusing on circuits (Elhage et al., 2021).
 A circuit represents a sparse oriented subgraph of a neural network having an interpretable functional role, where nodes correspond to computational units (e.g.
-neuron, attention head, RS hidden state) and edges represent computation paths Wang et al.
-(2023); Räuker et al.
-(2023).
+neuron, attention head, RS hidden state) and edges represent computation paths (Wang et al., 2023; Räuker et al., 2023).
 Formally, a circuit can be defined as $\mathcal{C}=(\mathcal{C}_{V},\mathcal{C}_{E})$ where $\mathcal{C}_{V}$ represents the set of computational nodes and $\mathcal{C}_{E}$ represents the set of directed edges between nodes where the information flows.
@@
 Task Specific Recurrent Circuits.
-A vast body of research has manually investigated task-specific circuits in Transformer-based LLM Bereska and Gavves (2024), such as entity resolution Wang et al.
-(2023), induction heads Olsson et al.
-(2022) or multi-hop reasoning Biran et al.
-(2024).
+A vast body of research has manually investigated task-specific circuits in Transformer-based LLM (Bereska and Gavves, 2024), such as entity resolution (Wang et al., 2023), induction heads (Olsson et al., 2022) or multi-hop reasoning (Biran et al., 2024).
 Analyzing the faithfulness of an NLE for such tasks would consists in directly drawing inspiration from these established circuits to define $\mathcal{C}$.
 Automated Circuit Discovery.
-Alternatively, circuits can be obtained through automated methods, including techniques based on activation patching Meng et al.
-(2022); Syed et al.
-(2024); Conmy et al.
-(2023), backward attribution Ferrando et al.
-(2022); Ferrando and Voita (2024) or Transcoders Dunefsky et al.
-(2024); Ameisen et al.
-(2025).
+Alternatively, circuits can be obtained through automated methods, including techniques based on activation patching (Meng et al., 2022; Syed et al., 2024; Conmy et al., 2023), backward attribution (Ferrando et al., 2022; Ferrando and Voita, 2024) or Transcoders (Dunefsky et al., 2024; Ameisen et al., 2025).
 These automated approaches help address scenarios where circuits may not be well-documented in the literature.
@@
 Concept-based Interpretation.
-Recent research has shown that interpretable high-level features, referred to as concepts, appear to be linearly encoded within LLM latent space Elhage et al.
-(2022); Park et al.
-(2024).
-This computational understanding of concepts aligns with definitions from cognitive science Ruiz Luyten and van der Schaar (2024), where concepts are considered as mental entities essential to thought, enabling information integration and categorization Goguen (2005).
+Recent research has shown that interpretable high-level features, referred to as concepts, appear to be linearly encoded within LLM latent space (Elhage et al., 2022; Park et al., 2024).
+This computational understanding of concepts aligns with definitions from cognitive science (Ruiz Luyten and van der Schaar, 2024), where concepts are considered as mental entities essential to thought, enabling information integration and categorization (Goguen, 2005).
 This alignment makes concepts an ideal granularity level for model interpretation.
@@
 We make the general assumption that no such set of labeled concepts is available.
-Generating concept-based interpretations can be done by applying dictionary learning techniques Tošić and Frossard (2011) to project hidden states into a concept space.
+Generating concept-based interpretations can be done by applying dictionary learning techniques (Tošić and Frossard, 2011) to project hidden states into a concept space.
 The semantics of the concepts are then determined once the concept dictionary has been learned.
-A popular approach for learning such dictionaries is to train sparse autoencoders (SAEHuben et al.
-(2023); Bricken et al.
-(2023) on the hidden states of $f$.
+A popular approach for learning such dictionaries is to train sparse autoencoders (SAE(Huben et al., 2023; Bricken et al., 2023) on the hidden states of $f$.
 Sparse concept activations are retrieved in the SAE hidden space to construct the concept-based activations as an interpretation.
@@
 Another way of interpreting $f$ hidden states is to generate free text interpretations.
-Two recent and similar approaches, namely Selfie Chen et al.
-(2024) and Patchscopes Ghandeharioun et al.
-(2024), perform such interpretation generation by leveraging LLM decoding capabilities to produce natural language descriptions of hidden states.
+Two recent and similar approaches, namely Selfie (Chen et al., 2024) and Patchscopes (Ghandeharioun et al., 2024), perform such interpretation generation by leveraging LLM decoding capabilities to produce natural language descriptions of hidden states.
 The interpretation is done by replacing the embedding of a placeholder token during a forward pass with the hidden state to be interpreted.
@@
 Local Faithfulness.
-For $\mathcal{I}=\texttt{SAE}$, each interpretation in $i(x)$ is a set of vectors of concept activations with dimensions potentially exceeding a million Templeton et al.
-(2024), making direct comparison challenging.
+For $\mathcal{I}=\texttt{SAE}$, each interpretation in $i(x)$ is a set of vectors of concept activations with dimensions potentially exceeding a million (Templeton et al., 2024), making direct comparison challenging.
 For a given interpretation $i^{\ell}_{k}(x)$, one way to simplify this comparison is to retrieve the concept labels with the highest activations.
@@
 In this case, an interpretation $i^{\ell}_{k}(x)$ can be compared to $e(x)$ by assessing whether one logically corroborates the other.
-This comparison can be done either under an LLM-as-a-Judge settings Gu et al.
-(2024) or through human evaluation.
+This comparison can be done either under an LLM-as-a-Judge settings (Gu et al., 2024) or through human evaluation.
 We give a detailed example of how this comparison can be carried out in the case of 2-hop reasoning in Section 4.
@@
 4.1 Task Description
-Multi-hop reasoning is a complex cognitive task in NLP that requires connecting a sequence of information to reach a conclusion Mavi et al.
-(2024).
-It consists in several single-hop operations Trivedi et al.
-(2022), which can individually be defined as triplets $(o_{1},r,o_{2})$ where $o_{1}$ is a source object, $r$ is a relation and $o_{2}$ is a target object.
+Multi-hop reasoning is a complex cognitive task in NLP that requires connecting a sequence of information to reach a conclusion (Mavi et al., 2024).
+It consists in several single-hop operations (Trivedi et al., 2022), which can individually be defined as triplets $(o_{1},r,o_{2})$ where $o_{1}$ is a source object, $r$ is a relation and $o_{2}$ is a target object.
 For example, the 2-hop reasoning statement "The country of origin of the movie maker that directed the movie Persona is Sweden" requires sequentially solving the two single-hop operations: $(o_{1},r_{1},o_{2})$ and $(o_{2},r_{2},o_{3})$ where $o_{1}$ is "Persona" (the movie), $r_{1}$ is "movie direction", $o_{2}$ is "Ingmar Bergman", $r_{2}$ is "country of origin" and the final answer $o_{3}$ is "Sweden".
@@
 Definition 5: Latent 2-hop Correctness If $\bm{\widehat{o_{3}}=o_{3}}$ and if $\exists(k,l)\in\mathcal{C_{V}}$ such that $o_{2}\in i^{\ell}_{k}(x)$, $f$ correctly performs latent 2-hop reasoning.
-This characterization of latent 2-hop reasoning is consistent with the one already mentioned in previous work Yang et al. (2024); Biran et al. (2024).
+This characterization of latent 2-hop reasoning is consistent with the one already mentioned in previous work (Yang et al., 2024; Biran et al., 2024).
 4.4 Highlighting the Tension Between Correctness and Faithfulness of 2-hop Self-NLE
@@
   • ($C_{4}$) $f$ being deceptive.
-If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$, $\widehat{o_{2}}\neq o_{2}$ and $o_{2}\in i^{\ell}_{k}(x)$, $f$ appears to correctly perform latent reasoning but does not faithfully express this in its self-NLE, this way doing deceptive reasoning Greenblatt et al.
-(2024).
+If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$, $\widehat{o_{2}}\neq o_{2}$ and $o_{2}\in i^{\ell}_{k}(x)$, $f$ appears to correctly perform latent reasoning but does not faithfully express this in its self-NLE, this way doing deceptive reasoning (Greenblatt et al., 2024).
 This case is expected to be rare, otherwise highlighting a case where $f$ is not honest in its self-NLE while knowing the ground truth bridge object, raising a problem in $f$ alignment.
   • ($C_{5}$) $f$ doing shortcut learning.
-If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\mathcal{C_{V}},o_{2}\notin i^{\ell}_{k}(x)$, the prediction is correct while $f$ seems to neither correctly perform latent reasoning nor generate a correct and faithful self-NLE, suggesting a potential case of shortcut learning Geirhos et al.
-(2020).
+If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\mathcal{C_{V}},o_{2}\notin i^{\ell}_{k}(x)$, the prediction is correct while $f$ seems to neither correctly perform latent reasoning nor generate a correct and faithful self-NLE, suggesting a potential case of shortcut learning (Geirhos et al., 2020).
 This case is different from $C_{2}$ since shortcut learning implies correctly predicting $o_{3}$ directly based on $o_{1}$, whereas when $f$ is biased, the $o_{3}$ prediction is based on an incorrect bridge object $o_{2}$.
@@
 5.1 Datasets and Models
-We apply NeuroFaith on the Wikidata-2-hop dataset Biran et al.
-(2024).
+We apply NeuroFaith on the Wikidata-2-hop dataset (Biran et al., 2024).
 Following the notations introduced in Section 4, each input text is formatted as $x=(o_{1},r_{1},{\blacktriangle{}},r_{2},\bullet)$.
-We evaluate Gemma-2-2B and Gemma-2-9B Riviere et al.
-(2024) in their instruction-tuned versions, using the following backbones from the transformers library Wolf et al.
-(2020): gemma-2-2B-it and gemma-2-9B-it.
+We evaluate Gemma-2-2B and Gemma-2-9B (Riviere et al., 2024) in their instruction-tuned versions, using the following backbones from the transformers library (Wolf et al., 2020): gemma-2-2B-it and gemma-2-9B-it.
 5.2 Framework Instantiation
@@
     Hidden States Interpretation.
-We use Selfie Chen et al.
-(2024) as interpreter $\mathcal{I}$ to decode the $f$ hidden states of interest and generate a set of local neural interpretations $i(x)$.
-We focus on the residual stream ($\mathcal{L}=\texttt{RS}$) and define the circuit $\mathcal{C}$ based on recent work Biran et al.
-(2024); Yang et al.
-(2024), which identifies the latent first hop operation in the early layers (when processing the last source token corresponding $o_{1}$).
+We use Selfie (Chen et al., 2024) as interpreter $\mathcal{I}$ to decode the $f$ hidden states of interest and generate a set of local neural interpretations $i(x)$.
+We focus on the residual stream ($\mathcal{L}=\texttt{RS}$) and define the circuit $\mathcal{C}$ based on recent work (Biran et al., 2024; Yang et al., 2024), which identifies the latent first hop operation in the early layers (when processing the last source token corresponding $o_{1}$).
 We therefore set the index token $k$ as the last one related to $o_{1}$ and $\ell\in\left\{{5,6,7,8,9,10,11}\right\}$ for gemma-2-2B-it and $\ell\in\left\{{8,9,10,11,12,13,14}\right\}$ for gemma-2-9B-it.
@@
 Our current study calculated faithfulness of free text self-NLE only in predict-then-explain contexts.
-A promising extension would be to compare the faithfulness of self-NLE with the performance gains they can induce Bhan et al.
-(2024).
-We believe that NeuroFaith could also lead to valuable results by measuring CoT faithfulness and being compared to other studies in this direction Lanham et al.
-(2023); Turpin et al.
-(2023).
-Examining the faithfulness of alternative explanation methods such as counterfactuals Madsen et al.
-(2024) or concept-based approaches Bhan et al.
-(2025) through comparison with the model’s internal neural activity is another promising research direction.
+A promising extension would be to compare the faithfulness of self-NLE with the performance gains they can induce (Bhan et al., 2024).
+We believe that NeuroFaith could also lead to valuable results by measuring CoT faithfulness and being compared to other studies in this direction (Lanham et al., 2023; Turpin et al., 2023).
+Examining the faithfulness of alternative explanation methods such as counterfactuals (Madsen et al., 2024) or concept-based approaches (Bhan et al., 2025) through comparison with the model’s internal neural activity is another promising research direction.
 7 Conclusion